# Repeating partial disruptions and two-body relaxation

### v1

@@ 4 Discussion and conclusions
We presented a simple semi-analytical model of partial disruptions for two stellar models, one for main sequence stars with mass $1\,M_{\odot}$ and the other for main sequence stars with mass $0.3\,M_{\odot}$. The prescriptions included in our semi-analytical models are based on the results of hydrodynamical simulations and semi-analytic modeling (Ryu et al.(2020a)Ryu, Krolik, Piran, & Noble; Ryu et al.(2020b)Ryu, Krolik, Piran, & Noble; Ryu et al.(2020c)Ryu, Krolik, Piran, & Noble; Ryu et al.(2020d)Ryu, Krolik, Piran, & Noble; Guillochon & Ramirez-Ruiz(2013); Gafton et al.(2015)Gafton, Tejeda, Guillochon, Korobkin, & Rosswog; Kremer et al.(2022)Kremer, Lombardi, Lu, Piro, & Rasio; Stone et al.(2013)Stone, Sari, & Loeb; Ivanov & Novikov(2001)). To properly study the phenomenon in the empty loss cone regime, where relaxation needs more than one orbit to change the orbital parameters, we accounted for the possibility of repeating partial disruptions. We will now discuss the implications of our model for the inclusion of partial disruptions in the loss cone theory, and the computation of their event rate.
@@ 4.1 Partial disruptions and loss cone theory
In the theory of the loss cone, one can use a Monte Carlo or a Fokker Planck approach to study the evolution of the distribution function of a stellar system (Cohn & Kulsrud(1978); Shapiro & Marchant(1978)). These studies show that particles enter the critical radius of the loss cone once are driven there by the stochastic encounters; in particular they typically start with a pericentre larger than the critical value and are then driven on orbits penetrating the critical sphere.
In the full loss cone regime, particles approaching the loss cone boundary can be scattered at any pericentre value (Merritt(2013)), so that all possible outcomes (i.e. partial disruption with orbital and mass change, ejection or total disruption) can occur. Neglecting the change in energy and mass, one can apply the classical loss cone theory - based on the instantaneous disruption model - to account for ejections and therefore use the corresponding (larger) critical pericentre.
@@
Stars described by model B, on the other hand, show a difference: the specific binding energy is subject to significant fluctuations, as these compact stars are more susceptible to tidal excitation. The orbital energy can change by more than $10\%$ as the star approaches the loss cone boundary, with a bias towards tighter configurations when considering Kremer’s estimate for mass loss (Eq. 10). This means that the star may be moved to regions where $q$ is smaller and two-body interactions less efficient. This process is somehow similar to that of extreme mass ratio inspirals (EMRIs), a phenomenon where a compact object approaches the central MBH so close that its evolution is deterministically driven by GWs emission over hundreds of thousand of orbits (Amaro-Seoane(2018)). In this case, the role of the emission of GWs at each pericentre passage would be replaced by a small mass loss and consequent tidal excitation. The total event rate of EMRIs (but not the distribution at capture) is computed by integrating the classically expected rate – with the loss cone radius equal to the innermost parabolic orbit – over the region where one expects EMRIs to be formed (Hopman & Alexander(2005); Bar-Or & Alexander(2016)). For stars of model B, a tentative approach to compute the total number of stars undergoing a partial disruption may be therefore to integrate the classical rate over the phase-space region where $N_{q}\gtrsim 2$. However, the information about the number of repeated partial disruptions that each of these stars undergoes must be inserted with a more detailed model.
As we mentioned at the beginning of this section, relaxation in the empty loss cone regime mainly drives particles from larger to smaller angular momenta/pericentres. This means that stars will gradually approach the critical thresholds identified in Fig. 5 from right to left in the plots, effectively resulting in a sequence of weak partial disruptions as shown in Fig. 6. As $q$ decreases, the mass lost at each pericentre passage is smaller, and therefore the resulting electromagnetic flare is expected to be weaker (Stone et al.(2020)Stone, Vasiliev, Kesden, Rossi, Perets, & Amaro-Seoane; Rossi et al.(2021)Rossi, Stone, Law-Smith, Macleod, Lodato, Dai, & Mandel).
@@ 4.2 Implications for TDE rates
Currently available estimates of time-evolving TDE rates rely on the assumption that the relaxed profile in angular momentum of stars is described by the CK solution, either in the whole range of $\mathcal{R}$, as in 1D Fokker Planck approaches (e.g. Vasiliev(2017); Stone & Metzger(2016)) or as the boundary condition at the loss cone interface for 2D models (e.g. Broggi et al.(2022)Broggi, Bortolas, Bonetti, Sesana, & Dotti). It is of primary interest to understand how these results should be interpreted when relaxing the hypothesis of instantaneous disruption.
@@
The efficiency of two-body relaxation in altering the orbital parameters when $q\gtrsim 1$ implies that the classically computed rates are reliable in this regime, and should be computed with a loss cone radius slightly larger than $\eta_{\mathrm{TTDE}}\,r_{\mathrm{TDE}}$ accounting for the possibility of ejections due to mass loss ($r_{\mathrm{LC}}^{A}\simeq 0.8\,r_{\mathrm{TDE}}$ and $r_{\mathrm{LC}}^{B}\simeq 1.2\,r_{\mathrm{TDE}}$ for the models we considered). Among the events computed this way, one should identify the pericentres resulting into observable events, as discussed in Bortolas et al.(2023)Bortolas, Ryu, Broggi, & Sesana.
@@
The case of partial disruptions in the empty loss cone regime instead requires a more careful treatment. As a first order effect, when $q=0.001$, particles at pericentres up to $r_{\mathrm{p}}\simeq 4.5\,r_{\mathrm{TDE}}$ for model A and $r_{\mathrm{p}}\simeq 2.05$ for model B will be affected by the presence of the loss cone. At smaller value of $q$, therefore higher specific binding energy, we expect this pericentre to increase due to the higher number of orbits allowed. This will correspond to a larger rate of partial disruptions, possibly repeating. Moreover, both models predict a significant change of the specific binding energy for particles as they approach the identified total disruption zone, with higher values of $q$ corresponding to stronger alterations. In the case of model A stars, part of the classically predicted rate will therefore result in failed disruptions, since particles will be moved to higher $q$ areas as the sequence of weak partial disruptions takes place. On the other hand, model B stars might increase their specific binding energy completing the sequence of weak disruptions. In both cases, the inclusion of partial disruptions in Fokker-Planck solvers requires a proper modeling of the phenomenon accounting for modifications to the Cohn-Kulsrud profile, see Eq. 18.
@@
For all models, the inclusion of repeated partial disruptions implies that a fraction of the classically expected TDEs will actually correspond to a sequence of weak disruptions, that may be too weak to be detected (Rossi et al.(2021)Rossi, Stone, Law-Smith, Macleod, Lodato, Dai, & Mandel; Stone et al.(2020)Stone, Vasiliev, Kesden, Rossi, Perets, & Amaro-Seoane), but can still contribute to the growth of the MBH. The large number of consecutive desruptions with weak mass loss may result in a sequence of bursts of gravitational waves, possibly increasing the gravitational wave background from TDEs (Toscani et al.(2020)Toscani, Rossi, & Lodato).
@@
It has been estimated that most of the TDEs originates in the empty loss cone regime, with a fraction approximately given by (Stone & Metzger(2016), eq. 29)
@@
corresponding to $70\%$ of the total expected cosmological rate. A qualitative result of our analysis is that part of these classically expected events will correspond to non-detectable pTDEs, therefore possibly reconciling the observed discrepancy between observed and expected TDE rates (Stone et al.(2020)Stone, Vasiliev, Kesden, Rossi, Perets, & Amaro-Seoane; van Velzen(2018)), while mantaining the possibility that stellar captures contribute to the growth of MBHs.
@@ 4.3 Caveats
In the case of model A, we determined the mass lost at the pericentre passage through Eq. 5. This equation has been derived for a MESA star with the core hydrogen mass fraction of 0.5 (Ryu et al.(2020a)Ryu, Krolik, Piran, & Noble) on a parabolic orbit. On the other hand, the effect of tidal excitation has been included through a prescription valid for $n=3.0$ polytropes. The MESA profile and the polytropic model correspond to different tidal excitation, with MESA model subject to stronger oscillations as shown in Generozov et al.(2018)Generozov, Stone, Metzger, & Ostriker (the corresponding function $T(\beta)$ is larger by a factor of a few).
@@
Liu et al.(2024)Liu, Ryu, Goodwin, Rau, Homan, Krumpe, Merloni, Grotova, Anderson, Malyali, & Miller-Jones simulated the partial disruption of different stars on orbits with eccentricity $e\simeq 0.95$ around an MBH with mass $M_{\bullet}=10^{5}\,M_{\odot}$. The simulations include the disruption of a $1\,M_{\odot}$ star with radius $1R_{\odot}$ initially described by a MESA model (corresponding to our model A stars). At small penetration factors, corresponding to weak mass loss, the simulated stars get slightly more tightly bound to the central MBH, the opposite of our predictions, showing a trend similar to what we expect for the less concentrated model B stars. In addition to the inconsistency between the MESA profile and the polytrope, this qualitative difference may also arise from the specific orbital parameters. In fact Eq.5 (the function relating the pericentre distance and the mass lost at the pericentre passage 5) has been fit to the simulation of disruptions on $e\gtrsim 0.99$ orbits (Ryu et al.(2020a)Ryu, Krolik, Piran, & Noble; Ryu et al.(2020b)Ryu, Krolik, Piran, & Noble; Ryu et al.(2020c)Ryu, Krolik, Piran, & Noble), and the tidal excitation prescription similarly assume a parabolic orbit (Lee & Ostriker(1986); Ivanov & Novikov(2001)). In general, the time spent at the pericentre increases as the eccentricity decreases, affecting both asymmetric mass loss and tidal excitation.
@@
The model we built for partial disruptions holds only on a range of MBH masses. Repeated disruptions at small penetration factors may completely disrupt a star as long as the motion of the central MBHs around the centre of the system is negligible with respect to the orbital timescales, setting a lower bound to the mass of the central MBH. On the other hand, the star can undergo a disruption as long as it is not gravitationally captured before the disruption, setting an upper bound to the mass of the central black hole ($M_{\bullet}\lesssim 10^{8}M_{\odot}$ for a sun-like star, the Hills mass). Finally, when the mass of the central black hole is large, one should account for general relativistic effects affecting the disruption process (Coughlin & Nixon(2022)) and the tidal excitation of the surviving core (Ivanov et al.(2003)Ivanov, Chernyakova, & Novikov).
Finally, we included relaxation in our model only by setting the maximum number of pericentre passages that a repeating pTDE can perform. However, a more complete treatment should include discrete orbital fluctuations (in energy and angular momentum), e.g. in a Monte Carlo fashion (Shapiro & Marchant(1978)). This will result, for example, in the possibility of interrupting the sequence of weak partial disruptions with a stronger event even in the empty loss cone regime (Weissbein & Sari(2017)).
@@ 4.4 Summary and Conclusions
We have constructed a simple model for the partial disruption of a star around a massive black hole. We estimated of the number of consecutive partial disruptions that a star will undergo depending on the strength of two-body encounters, and estimated the final outcome of the sequence of disruptions as a function of the initial pericentre distance. We considered two stellar models of $m_{\star}=1\,M_{\odot}$ and $m_{\star}=0.3\,M_{\odot}$, taking these two values as roughly representative of lower-main sequence and upper-main sequence internal structures. We then built a model for partial disruptions that accounts for the effects of mass loss and tidal excitation of the stellar structure. We built our model based on the results of hydrodynamical simulations (Ryu et al.(2020a)Ryu, Krolik, Piran, & Noble; Ryu et al.(2020b)Ryu, Krolik, Piran, & Noble; Ryu et al.(2020c)Ryu, Krolik, Piran, & Noble; Ryu et al.(2020d)Ryu, Krolik, Piran, & Noble; Guillochon & Ramirez-Ruiz(2013); Gafton et al.(2015)Gafton, Tejeda, Guillochon, Korobkin, & Rosswog) and previous theoretical studies (Kremer et al.(2022)Kremer, Lombardi, Lu, Piro, & Rasio; Stone et al.(2013)Stone, Sari, & Loeb; Ivanov & Novikov(2001)). For each model we studied the problem of single partial disruptions, arguing that this is the relevant problem for loss cone orbits ($r_{\mathrm{p}}$ comparable to $r_{\mathrm{TDE}}$) with binding energy small enough that two-body relaxation is efficient, i.e. in the full loss cone regime. We then considered the phenomenon of repeated partial disruptions, estimating the number of passages as the inverse of the loss cone occupation fraction $q$, as in Eq. (21). Therefore, pTDEs are expected to occur in the empty loss cone regime.
@@
  • Partial disruptions may alter the orbital energy without destroying the stars. For stars with $m_{\star}=1\,M_{\odot}$, weak pTDEs happening at pericentres larger than $\sim 1.5\,r_{\mathrm{TDE}}$ will result in a less tightly bound remnant. For stars with $m_{\star}=0.3\,M_{\odot}$, weak pTDEs happening at pericentres larger than $\sim 1.6\,r_{\mathrm{TDE}}$ will result in a more tightly bound remnan.
@@ Acknowledgements
LB thanks Alessandro Lupi and Martina Toscani for useful discussions. MB acknowledges support provided by MUR under grant “PNRR - Missione 4 Istruzione e Ricerca - Componente 2 Dalla Ricerca all’Impresa - Investimento 1.2 Finanziamento di progetti presentati da giovani ricercatori ID:SOE_0163” and by University of Milano-Bicocca under grant “2022-NAZ-0482/B”. LB, EB and AS acknowledge the financial support provided under the European Union’s H2020 ERC Consolidator Grant “Binary Massive Black Hole Astrophysics” (B Massive, Grant Agreement: 818691). EB acknowledges support from the European Union’s Horizon Europe programme under the Marie Skłodowska-Curie grant agreement No 101105915 (TESIFA), from the European Consortium for Astroparticle Theory in the form of an Exchange Travel Grant, and the European Union’s Horizon 2020 Programme under the AHEAD2020 project (grant agreement 871158). NCS was supported by the Israel Science Foundation (Individual Research Grant 2565/19) and the Binational Science Foundation (grant Nos. 2019772 and 2020397).
Software: the Julia language (Bezanson et al.(2017)Bezanson, Edelman, Karpinski, & Shah); Makie.jl (Danisch & Krumbiegel(2021)); Roots.jl (Verzani(2020)); Interpolations.jl (Kittisopikul & Holy(2023)).
Figure 7: Same as 6 but with the inclusion of the conversion of part of the oscillation energy $\epsilon_{\mathrm{osc}}$ into internal energy of the star through an enlargement of the stellar radius.

### v2

@@ 4 Discussion and conclusions
We presented a simple semi-analytical model of partial disruptions for two stellar models, one for main sequence stars with mass $1\,M_{\odot}$ and the other for main sequence stars with mass $0.3\,M_{\odot}$. The prescriptions included in our semi-analytical models are based on the results of hydrodynamical simulations and semi-analytic modeling (Ryu et al., 2020a, b, c, d; Guillochon & Ramirez-Ruiz, 2013; Gafton et al., 2015; Kremer et al., 2022; Stone et al., 2013; Ivanov & Novikov, 2001). To properly study the phenomenon in the empty loss cone regime, where relaxation needs more than one orbit to change the orbital parameters, we accounted for the possibility of repeating partial disruptions. We will now discuss the implications of our model for the inclusion of partial disruptions in the loss cone theory, and the computation of their event rate.
@@ 4.1 Partial disruptions and loss cone theory
In the theory of the loss cone, one can use a Monte Carlo or a Fokker Planck approach to study the evolution of the distribution function of a stellar system (Cohn & Kulsrud, 1978; Shapiro & Marchant, 1978). These studies show that particles enter the critical radius of the loss cone once are driven there by the stochastic encounters; in particular they typically start with a pericentre larger than the critical value and are then driven on orbits penetrating the critical sphere.
In the full loss cone regime, particles approaching the loss cone boundary can be scattered at any pericentre value (Merritt, 2013), so that all possible outcomes (i.e. partial disruption with orbital and mass change, ejection or total disruption) can occur. Neglecting the change in energy and mass, one can apply the classical loss cone theory - based on the instantaneous disruption model - to account for both ejections and disruptions. The standard treatment should be applied with the replacement of the tidal disruption radius $r_{\mathrm{TDE}}$ with the largest pericentre where ejections appear (as a function of energy). The critical pericentre corresponding to ejections is generally larger than $r_{\mathrm{TDE}}$, as can be seen in Fig. 5.
@@
Stars described by model B, on the other hand, show a difference: the specific binding energy is subject to significant fluctuations, as these compact stars are more susceptible to tidal excitation. The orbital energy can change by more than $10\%$ as the star approaches the loss cone boundary, with a bias towards tighter configurations when considering Kremer’s estimate for mass loss (Eq. 10). This means that the star may be moved to regions where $q$ is smaller and two-body interactions less efficient. This process is somehow similar to that of extreme mass ratio inspirals (EMRIs), a phenomenon where a compact object approaches the central MBH so close that its evolution is deterministically driven by GWs emission over hundreds of thousand of orbits (Amaro-Seoane, 2018). In this case, the role of the emission of GWs at each pericentre passage would be replaced by a small mass loss and consequent tidal excitation. The total event rate of EMRIs (but not the distribution at capture) is computed by integrating the classically expected rate – with the loss cone radius equal to the innermost parabolic orbit – over the region where one expects EMRIs to be formed (Hopman & Alexander, 2005; Bar-Or & Alexander, 2016). For stars of model B, a tentative approach to compute the total number of stars undergoing a partial disruption may be therefore to integrate the classical rate over the phase-space region where $N_{q}\gtrsim 2$. However, the information about the number of repeated partial disruptions that each of these stars undergoes must be inserted with a more detailed model.
As we mentioned at the beginning of this section, relaxation in the empty loss cone regime mainly drives particles from larger to smaller angular momenta/pericentres. This means that stars will gradually approach the critical thresholds identified in Fig. 6 from right to left in the plots, effectively resulting in a sequence of weak partial disruptions as shown in Fig. 7. As $q$ decreases, the mass lost at each pericentre passage is smaller, and therefore the resulting electromagnetic flare is expected to be weaker (Stone et al., 2020; Rossi et al., 2021).
@@ 4.2 Implications for TDE rates
Currently available estimates of time-evolving TDE rates rely on the assumption that the relaxed profile in angular momentum of stars is described by the CK solution, either in the whole range of $\mathcal{R}$, as in 1D Fokker Planck approaches (e.g. Vasiliev, 2017; Stone & Metzger, 2016) or as the boundary condition at the loss cone interface for 2D models (e.g. Broggi et al., 2022). It is of primary interest to understand how these results should be interpreted when relaxing the hypothesis of instantaneous disruption.
@@
The efficiency of two-body relaxation in altering the orbital parameters when $q\gtrsim 1$ implies that the classically computed rates are reliable in this regime, and should be computed with a loss cone radius slightly larger than $\eta_{\mathrm{TTDE}}\,r_{\mathrm{TDE}}$ accounting for the possibility of ejections due to mass loss ($r_{\mathrm{LC}}^{A}\simeq 0.8\,r_{\mathrm{TDE}}$ and $r_{\mathrm{LC}}^{B}\simeq 1.2\,r_{\mathrm{TDE}}$ for the models we considered). Among the events computed this way, one should identify the pericentres resulting into observable events, as discussed in Bortolas et al. (2023).
@@
The case of partial disruptions in the empty loss cone regime instead requires a more careful treatment. In the most extreme case we considered, $q=0.001$, stars with pericentres up to $r_{\mathrm{p}}\simeq 4.5\,r_{\mathrm{TDE}}$ for model A and $r_{\mathrm{p}}\simeq 2.05$ for model B will be affected by the presence of the loss cone. The trend emerging from our analysis is that at smaller value of $q$, therefore higher specific binding energy, the critical value of the pericentre, up to which orbits result in pTDEs, will increase due to the larger number of orbits allowed. This will correspond to a larger rate of partial disruptions, possibly repeating. Moreover, both models predict a significant change of the specific binding energy for particles as they approach the identified total disruption zone, with higher values of $q$ corresponding to stronger alterations. In the case of model A stars, part of the classically predicted rate will therefore result in failed disruptions, since particles will be moved to higher $q$ areas as the sequence of weak partial disruptions takes place. On the other hand, model B stars might increase their specific binding energy completing the sequence of weak disruptions. In both cases, the inclusion of partial disruptions in Fokker-Planck solvers requires a proper modelling of the phenomenon accounting for modifications to the Cohn-Kulsrud profile, see Eq. 18.
@@
For all models, the inclusion of repeated partial disruptions implies that a fraction of the classically expected TDEs will actually correspond to a sequence of weak disruptions, that may be too weak to be detected (Rossi et al., 2021; Stone et al., 2020), but can still contribute to the growth of the MBH. The large number of consecutive desruptions with weak mass loss may result in a sequence of bursts of gravitational waves, possibly increasing the gravitational wave background from TDEs (Toscani et al., 2020).
@@
It has been estimated that most of the TDEs originates in the empty loss cone regime, with a fraction approximately given by (Stone & Metzger, 2016, Eq. 29)
@@
corresponding to $70\%$ of the total expected cosmological rate. A qualitative result of our analysis is that part of these classically expected events will correspond to non-detectable pTDEs, therefore possibly reconciling the observed discrepancy between observed and expected TDE rates (Stone et al., 2020; van Velzen, 2018), while maintaining the possibility that stellar captures contribute to the growth of MBHs.
@@ 4.3 Caveats
In the case of model A, we determined the mass lost at the pericentre passage through Eq. 5. This equation has been derived for a MESA star with the core hydrogen mass fraction of 0.5 (Ryu et al., 2020a) on a parabolic orbit. On the other hand, the effect of tidal excitation has been included through a prescription valid for $n=3.0$ polytropes. The MESA profile and the polytropic model correspond to different tidal excitation, with MESA model subject to stronger oscillations as shown in Generozov et al. (2018) (the corresponding function $T(\beta)$ is larger by a factor of a few).
@@
Liu et al. (2024) simulated the partial disruption of different stars on orbits with eccentricity $e\simeq 0.95$ around an MBH with mass $M_{\bullet}=10^{5}\,M_{\odot}$. The simulations include the disruption of a $1\,M_{\odot}$ star with radius $1R_{\odot}$ initially described by a MESA model (corresponding to our model A stars). At small penetration factors, corresponding to weak mass loss, the simulated stars get slightly more tightly bound to the central MBH, the opposite of our predictions, showing a trend similar to what we expect for the less concentrated model B stars. In addition to the inconsistency between the MESA profile and the polytrope, this qualitative difference may also arise from the specific orbital parameters. In fact Eq.5 (the function relating the pericentre distance and the mass lost at the pericentre passage) has been fit to the simulation of disruptions on $e\gtrsim 0.99$ orbits (Ryu et al., 2020a, b, c), and the tidal excitation prescription similarly assume a parabolic orbit (Lee & Ostriker, 1986; Ivanov & Novikov, 2001). In general, the time spent at the pericentre increases as the eccentricity decreases, affecting both asymmetric mass loss and tidal excitation.
@@
The model we built for partial disruptions holds only on a range of MBH masses. Repeated disruptions at small penetration factors may completely disrupt a star as long as the motion of the central MBHs around the centre of the system is negligible with respect to the orbital timescales, setting a lower bound to the mass of the central MBH. On the other hand, the star can undergo a disruption as long as it is not gravitationally captured before the disruption, setting an upper bound to the mass of the central black hole ($M_{\bullet}\lesssim 10^{8}M_{\odot}$ for a sun-like star, the Hills mass). Finally, when the mass of the central black hole is large, one should account for general relativistic effects affecting the disruption process (Coughlin & Nixon, 2022) and the tidal excitation of the surviving core (Ivanov et al., 2003).
Finally, we included relaxation in our model only by setting the maximum number of pericentre passages that a repeating pTDE can perform. However, a more complete treatment should include discrete orbital fluctuations (in energy and angular momentum), e.g. in a Monte Carlo fashion (Shapiro & Marchant, 1978). This will result, for example, in the possibility of interrupting the sequence of weak partial disruptions with a stronger event even in the empty loss cone regime (Weissbein & Sari, 2017).
@@ 4.4 Summary and Conclusions
We have constructed a simple model for the partial disruption of a star around a massive black hole. We estimated of the number of consecutive partial disruptions that a star will undergo depending on the strength of two-body encounters, and estimated the final outcome of the sequence of disruptions as a function of the initial pericentre distance. We considered two stellar models of $m_{\star}=1\,M_{\odot}$ and $m_{\star}=0.3\,M_{\odot}$, taking these two values as roughly representative of lower-main sequence and upper-main sequence internal structures. We then built a model for partial disruptions that accounts for the effects of mass loss and tidal excitation of the stellar structure. We built our model based on the results of hydrodynamical simulations (Ryu et al., 2020a, b, c, d; Guillochon & Ramirez-Ruiz, 2013; Gafton et al., 2015) and previous theoretical studies (Kremer et al., 2022; Stone et al., 2013; Ivanov & Novikov, 2001). For each model we studied the problem of single partial disruptions, arguing that this is the relevant problem for loss cone orbits ($r_{\mathrm{p}}$ comparable to $r_{\mathrm{TDE}}$) with binding energy small enough that two-body relaxation is efficient, i.e. in the full loss cone regime. We then considered the phenomenon of repeated partial disruptions, estimating the number of passages as the inverse of the loss cone diffusivity $q$, as in Eq. (21). Therefore, pTDEs are expected to occur in the empty loss cone regime.
@@
  • Partial disruptions may alter the orbital energy without destroying the stars. For stars with $m_{\star}=1\,M_{\odot}$, the energy change is driven by mass loss, and weak pTDEs happening at pericentres larger than $\sim 1.5\,r_{\mathrm{TDE}}$ will result in a less tightly bound remnant. For stars with $m_{\star}=0.3\,M_{\odot}$, the energy change is driven by tidal oscillations, and weak pTDEs happening at pericentres larger than $\sim 1.6\,r_{\mathrm{TDE}}$ will result in a more tightly bound remnant at the first pericentre passage.
@@ Acknowledgements
LB thanks David Izquierdo Villalba, Alessandro Lupi, and Martina Toscani for useful discussions. We thank Julian Krolik for useful comments. We thank the anonymous referee for their useful comments and suggestions. MB acknowledges support provided by MUR under grant “PNRR - Missione 4 Istruzione e Ricerca - Componente 2 Dalla Ricerca all’Impresa - Investimento 1.2 Finanziamento di progetti presentati da giovani ricercatori ID:SOE_0163” and by University of Milano-Bicocca under grant “2022-NAZ-0482/B”. LB, EB and AS acknowledge the financial support provided under the European Union’s H2020 ERC Consolidator Grant “Binary Massive Black Hole Astrophysics” (B Massive, Grant Agreement: 818691). EB acknowledges support from the European Union’s Horizon Europe programme under the Marie Skłodowska-Curie grant agreement No 101105915 (TESIFA), from the European Consortium for Astroparticle Theory in the form of an Exchange Travel Grant, and the European Union’s Horizon 2020 Programme under the AHEAD2020 project (grant agreement 871158). NCS was supported by the Israel Science Foundation (Individual Research Grant 2565/19) and the Binational Science Foundation (grant Nos. 2019772 and 2020397).
Software: the Julia language (Bezanson et al., 2017); Makie.jl (Danisch & Krumbiegel, 2021); Roots.jl (Verzani, 2020); Interpolations.jl (Kittisopikul & Holy, 2023).
Figure 8: Same as 7 but with the inclusion of the conversion of part of the oscillation energy $\epsilon_{\mathrm{osc}}$ into internal energy of the star through an enlargement of the stellar radius.